OpenStudy (anonymous):

solve integratn 1/a+x^2+y^2 dy n plzz guide the method to solve it

OpenStudy (richyw):

can you please type this out in latex? just put \(\text{\[}\) before the equation and \(\text{\]}\) after the equation. Integral is \int. integral from a to b is int_a^b. You can use the equation editor as well if you don't know latex.

OpenStudy (anonymous):

\[int \[1\div1+x^2+y^2 dy\]

OpenStudy (richyw):

uh, try again. are you trying to say \[\int\left(\frac{1}{a}+x^2+y^2\right)dy\]

OpenStudy (anonymous):

yes can u reply bkk??

OpenStudy (richyw):

yup, starting from the equation above... \[\int\frac{1}{a}dy+\int x^2 dy+\int y^2 dy\]\[\frac{1}{a}\int dy+x^2\int dy+\int y^2 dy\]then do these basic integrations,\[\frac{1}{a}y+x^2y+\frac{y^3}{3}\]and finally, factoring out the \(y\) we have,\[y\left(\frac{1}{a}+x^2+\frac{y^2}{3}\right)\]